### v2

@@ Appendix A Model fitting results
A binomial model with a logit function of the positive response rate. Samples are generated using 4 chains with 2,500 post-warmup iterations each.
Fixed Effects
Estimate Est.Error l-95% CI u-95% CI R-hat Bulk_ESS Tail_ESS
@@
Standard Deviation of Varying Effects
Estimate Est.Error l-95% CI u-95% CI R-hat Bulk_ESS Tail_ESS
@@
Table A.1: Model A fit summaries.
Appendix B Model estimates
Figures B.1 and B.2 present the weekly incidence estimates by sex and age groups, respectively.
Figure B.1: Estimated weekly incidence by sex based on the multilevel regression and poststratification. Females tend to have lower infection rates than males with small weekly differences. The shaded areas represent 95% credible intervals.
Figure B.2: Estimated weekly incidence by age group based on the multilevel regression and poststratification. Young adults tend to have lower infection rates than elders with small differences during most weeks. The shaded areas represent 95% credible intervals.

### v3

@@ Appendix B Model fitting results
A binomial model with a logit function of the positive response rate. Samples are generated using 4 chains with 2500 post-warmup iterations each.
Non-varying Effects
Estimate Est.Error l-95% u-95% R-hat Bulk_ESS Tail_ESS
@@
Standard Deviations of Varying Effects
Estimate Est.Error l-95% u-95% R-hat Bulk_ESS Tail_ESS
@@
Table B.1: Model A fit summaries.
Appendix C Model estimates
Figures C.1 and C.2 present the weekly incidence estimates by sex and age groups, respectively.
Figure C.1: Estimated weekly incidence by sex based on the multilevel regression and poststratification. Females tend to have lower infection rates than males with small weekly differences. The shaded areas represent 95% credible intervals.
Figure C.2: Estimated weekly incidence by age group based on the multilevel regression and poststratification. Young adults tend to have lower infection rates than elders with small differences during most weeks. The shaded areas represent 95% credible intervals.